# How Data Mixing Shapes In-Context Learning: Asymptotic Equivalence for Transformers with MLPs

### v1

@@ Abstract
Pretrained Transformers demonstrate remarkable in-context learning (ICL) capabilities, enabling them to adapt to new tasks from demonstrations without parameter updates. However, theoretical studies often rely on simplified architectures (e.g., omitting MLPs), data models (e.g., linear regression with isotropic inputs), and single-source training—limiting their relevance to realistic settings. In this work, we study ICL in pretrained Transformers with nonlinear MLP heads on nonlinear tasks drawn from multiple data sources with heterogeneous input, task, and noise distributions. We analyze a model where the MLP comprises two layers, with the first layer trained via a single gradient step and the second layer fully optimized. Under high-dimensional asymptotics, we prove that such models are equivalent in ICL error to structured polynomial predictors, leveraging results from the theory of Gaussian universality and orthogonal polynomials. This equivalence reveals that nonlinear MLPs meaningfully enhance ICL performance—particularly on nonlinear tasks—compared to linear baselines. It also enables a precise analysis of data mixing effects: we identify key properties of high-quality data sources (low noise, structured covariances) and show that feature learning emerges only when the task covariance exhibits sufficient structure. These results are validated empirically across various activation functions, model sizes, and data distributions. Finally, we experiment with a real-world scenario involving multilingual sentiment analysis where each language is treated as a different source. Our experimental results for this case exemplify how our findings extend to real-world cases. Overall, our work advances the theoretical foundations of ICL in Transformers and provides actionable insight into the role of architecture and data in ICL.
@@ 1 Introduction
Transformers vaswani2017attention have emerged as a dominant architecture across a broad range of machine learning applications. A particularly striking capability of these models is in-context learning (ICL)—the ability to adapt to new tasks from a few examples presented as input, without parameter updates brown2020language . As this phenomenon becomes central to modern AI systems, a deeper theoretical understanding of when and why Transformers succeed at ICL is increasingly critical.
To make this challenge tractable, recent work has studied simplified settings, often focusing on linear tasks and attention-only models that omit multilayer perceptrons (MLPs) akyurek2023what ; pmlr-v202-von-oswald23a ; wu2024how ; zhang2024trained . However, such abstractions overlook the empirical observation that nonlinear MLPs are integral to Transformer performance in practice. While a few recent studies have begun to investigate MLPs in the ICL context kim2024transformers ; li2024nonlinear ; oko2024pretrained , their analyses are restricted either to nonstandard architectures, specific activation types, or overly simplistic data settings—typically assuming a single, homogeneous data source.
@@
To address these questions, we analyze the ICL performance of Transformers with a nonlinear MLP head trained under multiple data sources with different distributions, focusing on nonlinear regression tasks in high-dimensional regimes. We adopt an asymptotic lens where sample size, context length, and hidden dimension jointly diverge, and draw on Gaussian universality theory hu2022universality ; demir2025asymptotic to analytically characterize model behavior. Specifically, we study a Transformer with linear attention and a two-layer nonlinear MLP head, where the first layer is trained with a single gradient step and the second is fully trained. We show that this model is asymptotically equivalent to a finite-degree polynomial model in terms of ICL error. This equivalence allows us to (i) explain the performance gains due to nonlinear MLPs and (ii) study how different training data distributions affect ICL outcomes.
@@ 2 Related Work
The discovery of in-context learning (ICL) in large-scale Transformers brown2020language has prompted a surge of empirical and theoretical investigations. Empirically, studies such as wei2022emergent ; olsson2022context ; schaeffer2023are highlight that ICL capabilities grow with model size, suggesting an emergent behavior critical to modern foundation models. On the theoretical side, controlled synthetic tasks—especially linear regression—have become a standard benchmark for analyzing ICL mechanisms in simplified settings zhang2024trained ; garg2022can ; raventos2024pretraining . These studies often adopt linearized attention mechanisms for tractability and argue that Transformers learn to simulate algorithmic procedures during pretraining bai2023transformers ; akyurek2023what ; ahn2023transformers ; li2023transformers ; pmlr-v202-von-oswald23a ; mahankali2024one ; fu2024transformers ; li2024finegrained ; park2024competition . Yet, the precise nature of these learned procedures remains unclear. Notably, zhang2024trained ; lu2024incontext ; lu2025asymptotic offer generalization analysis for Transformers with a linear attention (without MLPs) in linear ICL tasks, laying the groundwork for extensions to more realistic settings.
@@
While much of the theoretical literature abstracts away the MLP components of Transformers, recent efforts have started to examine their impact on ICL. For example, li2024nonlinear studied ReLU-based MLPs in classification tasks, while kim2024transformers ; oko2024pretrained analyzed architectures where MLP layers precede attention—deviating from the standard Transformer design in which MLPs follow attention blocks. Furthermore, abedsoltan2024context studied a simplified Transformer architecture with a fixed (non-trainable) nonlinear attention mechanism—omitting the training of key, query, and value weights—and an optional MLP head. Overall, these studies either restrict task types, limit activation functions, adopt nonstandard architectures, or omit training of the attention weights, leaving open questions about the role of nonlinear MLPs in Transformers for nonlinear ICL in realistic settings. Our work addresses this gap by analyzing a Transformer with a trained nonlinear MLP head consistent with the canonical architecture and showing asymptotic equivalence of a polynomial surrogate model to the Transformer model in terms of the ICL error. Moreover, prior works focus primarily on single-source training with simplified data distributions, whereas we consider multiple heterogeneous data sources and explicitly characterize the impact of data mixture ratios on the ICL performance.
@@
The concept of Gaussian universality from random matrix theory has proven valuable for the asymptotic analysis of neural networks. Under isotropic Gaussian inputs, random feature models (two-layer neural network with random first layer and trained second layer) can be shown to be equivalent to Gaussian models with matched first and second moments goldt2022gaussian ; hu2022universality ; couillet2022random . These results have been extended to broader settings, including empirical risk minimization montanari2022universality , mixture-distributed inputs dandi2023universality , and structured covariances demir2024randomfeatures . More recently, the equivalence has been rigorously established for two-layer networks trained with one gradient step moniri2023theory ; dandi2023learning ; demir2025asymptotic . However, these universality results have not yet been extended to Transformer architectures or in-context learning scenarios. Our work bridges this gap by showing that a Transformer with a nonlinear MLP is asymptotically equivalent to a finite-degree polynomial model—thus connecting Gaussian universality theory to Transformer-based ICL for complex data scenarios involving data mixing.
@@ 3.2 Data model
While most of the theoretical works assume a single data source, in practice, the models are pretrained with datasets consisting of a mixture of data sources ye2025data . To reflect this, we consider $\mathcal{S}$ different data sources. Thus, when constructing a dataset, we first sample the source index $s$ from a categorical distribution with the following probabilities: $\mathbb{P}(s=i)=\rho_{i}$ for $\{0,1,\dots,\mathcal{S}-1\}$. Then, for each data source, we posit a nonlinear relationship between $\bm{x}$} and $y$ governed by a context-specific parameter vector $\bm{\xi}\in\mathbb{R}^{d}$, subject to additive Gaussian noise. Although $\bm{\xi}$ remains constant within a given context, it is re-sampled across different contexts, imposing the requirement that the model estimate $\bm{\xi}$ from the observed pairs before generalizing to the new input.
@@
Following zhang2024trained , we construct an embedding matrix ${\bm{Z}}\in\mathbb{R}^{(d+1)\times(\ell+1)}$ by stacking the feature vectors and their corresponding labels, using a zero placeholder for the unknown label:
@@ 3.3 Transformer model
In this work, we consider a Transformer model involving single block / layer of a linear attention and a MLP head (two-layer neural network). While it is possible to extend the setting to those involving multiple blocks of attention and MLP layers, following many theoretical studies zhang2024trained ; lu2024incontext ; akyurek2023what in the literature, we focus on the single block case to streamline our theoretical analysis. Similarly, while nonlinear attention mechanisms (such as softmax attention) are popular in the empirical literature, we use linear attention since it is analytically tractable zhang2024trained ; lu2024incontext and variants of linear attention can lead to performance similar to softmax attention han2024bridging . Another benefit of focusing on linear attention is that it allows us to isolate the role of the nonlinear MLP head in capturing the nonlinear nature of our ICL tasks since nonlinear attention can also play a role in capturing such nonlinear relations. In the rest of this section, we first describe the linear attention mechanism–introducing linear Transformer without MLP head as our baseline– and then, we explain our Transformer model with MLP head together with the training procedure.
@@ 3.3.1 Linear attention (linear Transformer)
where ${\bm{K}},{\bm{Q}},{\bm{V}}$ are appropriately sized key, query, and value matrices, respectively. When predicting $y_{\ell+1}$, the relevant output of the linear Transformer (without MLPs) will be $A_{d+1,\ell+1}$, i.e., the element of $\bm{A}$} corresponding to the 0 element in the embedding matrix $\bm{Z}$}. Using the reparameterization in lu2024incontext (see Appendix A), this prediction can be simplified to:
@@ 3.3.2 Transformer with a nonlinear MLP
To simplify our analysis while keeping feature learning for the MLP, we consider the following training procedure with two stages ba2022highdimensional ; damian2022neural :
@@
Here, by avoiding data reuse, we ensure $\hat{{\bm{F}}}$ remains independent of the second training set, which simplifies theoretical analysis ba2022highdimensional . Furthermore, while our training procedure differs from standard end-to-end training techniques, this two-phase scheme –gradient descent on the first layer followed by ridge regression on the second– is well-motivated by prior theoretical work on MLPs ba2022highdimensional ; damian2022neural . This separation enables tractable analysis while preserving meaningful feature learning dynamics.
@@ 3.4 Evaluating ICL Performance
where $\hat{y}$ denotes the model prediction, given by (4) for the linear Transformer or by (5) for the Transformer with a nonlinear MLP. This formulation enables controlled experimentation with training-time data mixtures while evaluating performance as an average across all sources. Such an approach aligns with practical multi-source training scenarios, where models are trained on heterogeneous datasets and validated by averaging performance over the constituent sources ye2025data . Note that, we also provide ICL errors corresponding to each source (for cases when it might be relevant) in Appendix G.
@@ 4.1 Assumptions
This assumption aligns with regimes studied in prior work: $\ell/d$ and $n/d^{2}$ are critical for ICL performance in linear Transformers lu2024incontext , while $k/n$ ensures the model capacity scales with data, consistent with the literature hu2022universality .
@@
The covariance matrices of input vectors satisfy $\|{\bm{\Sigma}}_{x,s}\|_{2}^{2}=\mathcal{O}(d)$ and $\text{Tr}({\bm{\Sigma}}_{x,s})\asymp d$ while the mean vectors satisfy $\|{\bm{\mu}}_{x,s}\|_{2}^{2}=\mathcal{O}(d)$.
@@
By permitting step size to scale with the input dimension while the dimension tends to infinity, we introduce non-trial feature learning, allowing the model to learn nonlinear features ba2022highdimensional ; demir2025asymptotic .
@@
$\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})|s=i))=\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})|s=j)),$
@@
$\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})|s)={\bm{I}}_{d(d+1)}+\sum_{q=1}^{r_{s}}\theta_{s,q}\bm{\gamma}_{s,q}\bm{\gamma}_{s,q}^{T},$ (9)
where $r_{s}\in\mathbb{Z}^{+}$, $\theta_{s,q}>0$, and $\{\bm{\gamma}_{s,q}\}$ are orthonormal vectors in $\mathbb{R}^{n}$.
@@
$\eta\cdot\|\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))\|=o(d^{2}).$
Assumptions 4.3- 4.5 ensure tractability in our asymptotic analysis while allowing the feature learning. Although these assumptions might look unnatural at first glance, they enable us to connect our setting involving ICL by Transformer with a nonlinear MLP to prior results demir2025asymptotic for MLPs, facilitating our theoretical analysis. We would like to note that Assumptions 4.3- 4.5 represent limitations of our theoretical results. Nevertheless, our experimental results on both synthetic and real-world datasets suggest that these assumptions can be partially relaxed in practice.
@@
The vector ${\bm{w}}\sim\mathcal{N}(0,\frac{{\bm{I}}_{k}}{k})$ and the matrix ${\bm{F}}:=[{\bm{f}}_{1},\dots,{\bm{f}}_{k}]^{T}$ has rows ${\bm{f}}_{i}\sim\mathcal{N}\left(0,\frac{{\bm{I}}_{n}}{\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))}\right)$.
This initialization is consistent with standard practices for MLP initializations goodfellow2016deep .
@@
The activation $\sigma:\mathbb{R}\to\mathbb{R}$ has bounded derivatives and satisfies $\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma(z)^{2}]<\infty$, admitting a Hermite expansion (O’Donnell_2014, , Chapter 11.2):

### v2

@@ Abstract
Pretrained Transformers demonstrate remarkable in-context learning (ICL) capabilities, enabling them to adapt to new tasks from demonstrations without parameter updates. However, theoretical studies often rely on simplified architectures (e.g., omitting MLPs), plain data models (e.g., linear regression with isotropic inputs), and single-source training—limiting their relevance to realistic settings. In this work, we study ICL in pretrained Transformers with nonlinear MLP heads on nonlinear tasks drawn from multiple data sources with heterogeneous input, task, and noise distributions. We analyze a model where the MLP comprises two layers, with the first layer trained via a single gradient step and the second layer fully optimized. Under high-dimensional asymptotics, we prove that such models are equivalent in ICL error to structured polynomial predictors, leveraging results from the theory of Gaussian universality and orthogonal polynomials. This equivalence reveals that nonlinear MLPs meaningfully enhance ICL performance—particularly on nonlinear tasks—compared to linear baselines. It also enables a precise analysis of data mixing effects: we identify key properties of high-quality data sources (low noise, structured covariances) and show that feature learning emerges only when the task covariance exhibits sufficient structure. These results are validated empirically across various activation functions, model sizes, and data distributions. Finally, we experiment with a real-world scenario involving multilingual sentiment analysis where each language is treated as a different source. Our experimental results for this case exemplify how our findings extend to real-world cases. Overall, our work advances the theoretical foundations of ICL in Transformers and provides actionable insight into the role of architecture and data in ICL.
@@ 1 Introduction
Transformers Vaswani et al. (2017) have emerged as a dominant architecture across a broad range of machine learning applications. A particularly striking capability of these models is in-context learning (ICL)—the ability to adapt to new tasks from a few examples presented as input, without parameter updates Brown et al. (2020). As this phenomenon becomes central to modern AI systems, a deeper theoretical understanding of when and why Transformers succeed at ICL is increasingly critical.
To make this challenge tractable, recent work has studied simplified settings, often focusing on linear tasks and attention-only models that omit multilayer perceptrons (MLPs) Akyürek et al. (2023); Von Oswald et al. (2023); Wu et al. (2024); Zhang et al. (2024). However, such abstractions overlook the empirical observation that nonlinear MLPs are integral to Transformer performance in practice. While a few recent studies have begun to investigate MLPs in the ICL context Kim and Suzuki (2024); Li et al. (2024a); Oko et al. (2024), their analyses are restricted either to nonstandard architectures, specific activation types, or overly simplistic data settings—typically assuming a single, homogeneous data source.
@@
To address these questions, we analyze the ICL performance of Transformers with a nonlinear MLP head trained under multiple data sources with different distributions, focusing on nonlinear regression tasks in high-dimensional regimes. We adopt an asymptotic lens where sample size, context length, and hidden dimension jointly diverge, and draw on Gaussian universality theory Hu and Lu (2023); Demir and Dogan (2025a) to analytically characterize model behavior. Specifically, we study a Transformer with linear attention and a two-layer nonlinear MLP head, where the first layer is trained with a single gradient step and the second is fully trained. We show that this model is asymptotically equivalent to a finite-degree polynomial model in terms of ICL error. This equivalence allows us to (i) explain the performance gains due to nonlinear MLPs and (ii) study how different training data distributions affect ICL outcomes.
@@ 2 Related Work
The discovery of in-context learning (ICL) in large-scale Transformers Brown et al. (2020) has prompted a surge of empirical and theoretical investigations. Empirically, studies such as Wei et al. (2022); Olsson et al. (2022); Schaeffer et al. (2023) highlight that ICL capabilities grow with model size, suggesting an emergent behavior critical to modern foundation models. On the theoretical side, controlled synthetic tasks—especially linear regression—have become a standard benchmark for analyzing ICL mechanisms in simplified settings Zhang et al. (2024); Garg et al. (2022); Raventós et al. (2024). These studies often adopt linearized attention mechanisms for tractability and argue that Transformers learn to simulate algorithmic procedures during pretraining Bai et al. (2023); Akyürek et al. (2023); Ahn et al. (2023); Li et al. (2023); Von Oswald et al. (2023); Mahankali et al. (2024); Fu et al. (2024); Li et al. (2024b); Park et al. (2024). Yet, the precise nature of these learned procedures remains unclear. Notably, Zhang et al. (2024); Lu et al. (2024, 2025) offer generalization analysis for Transformers with a linear attention (without MLPs) in linear ICL tasks, laying the groundwork for extensions to more realistic settings.
@@
While much of the theoretical literature abstracts away the MLP components of Transformers, recent efforts have started to examine their impact on ICL. For example, Li et al. (2024a) studied ReLU-based MLPs in classification tasks, while Kim and Suzuki (2024); Oko et al. (2024) analyzed architectures where MLP layers precede attention—deviating from the standard Transformer design in which MLPs follow attention blocks. Furthermore, Abedsoltan et al. (2024) studied a simplified Transformer architecture with a fixed (non-trainable) nonlinear attention mechanism—omitting the training of key, query, and value weights—and an optional MLP head. Overall, these studies either restrict task types, limit activation functions, adopt nonstandard architectures, or omit training of the attention weights, leaving open questions about the role of nonlinear MLPs in Transformers for nonlinear ICL in realistic settings. Our work addresses this gap by analyzing a Transformer with a trained nonlinear MLP head consistent with the canonical architecture and showing asymptotic equivalence of a polynomial surrogate model to the Transformer model in terms of the ICL error. Moreover, prior works focus primarily on single-source training with simplified data distributions, whereas we consider multiple heterogeneous data sources and explicitly characterize the impact of data mixture ratios on the ICL performance.
@@
The concept of Gaussian universality from random matrix theory has proven valuable for the asymptotic analysis of neural networks. Under isotropic Gaussian inputs, random feature models (two-layer neural network with random first layer and trained second layer) can be shown to be equivalent to Gaussian models with matched first and second moments Goldt et al. (2022); Hu and Lu (2023); Couillet and Liao (2022). These results have been extended to broader settings, including empirical risk minimization Montanari and Saeed (2022), mixture-distributed inputs Dandi et al. (2023b), and structured covariances Demir and Dogan (2024). More recently, the equivalence has been rigorously established for two-layer networks trained with one gradient step Moniri et al. (2024); Dandi et al. (2023a); Demir and Dogan (2025a). However, these universality results have not yet been extended to Transformer architectures or in-context learning scenarios. Our work bridges this gap by showing that a Transformer with a nonlinear MLP is asymptotically equivalent to a finite-degree polynomial model—thus connecting Gaussian universality theory to Transformer-based ICL for complex data scenarios involving data mixing.
@@ 3.2 Data model
While most of the theoretical works assume a single data source, in practice, the models are pretrained with datasets consisting of a mixture of data sources Ye et al. (2025). To reflect this, we consider $\mathcal{S}$ different data sources. Thus, when constructing a dataset, we first sample the source index $s$ from a categorical distribution with the following probabilities: $\mathbb{P}(s=i)=\rho_{i}$ for $\{0,1,\dots,\mathcal{S}-1\}$. Then, for each data source, we posit a nonlinear relationship between $\bm{x}$} and $y$ governed by a context-specific parameter vector $\bm{\xi}\in\mathbb{R}^{d}$, subject to additive Gaussian noise. Although $\bm{\xi}$ remains constant within a given context, it is re-sampled across different contexts, imposing the requirement that the model estimate $\bm{\xi}$ from the observed pairs before generalizing to the new input.
@@
Following Zhang et al. (2024), we construct an embedding matrix ${\bm{Z}}\in\mathbb{R}^{(d+1)\times(\ell+1)}$ by stacking the feature vectors and their corresponding labels, using a zero placeholder for the unknown label:
@@ 3.3 Transformer model
In this work, we consider a Transformer model involving single block / layer of a linear attention and an MLP head (two-layer neural network). While it is possible to extend the setting to those involving multiple blocks of attention and MLP layers, following many theoretical studies Zhang et al. (2024); Lu et al. (2024); Akyürek et al. (2023) in the literature, we focus on the single block case to streamline our theoretical analysis. Similarly, while nonlinear attention mechanisms (such as softmax attention) are popular in the empirical literature, we use linear attention since it is analytically tractable Zhang et al. (2024); Lu et al. (2024) and variants of linear attention can lead to performance similar to softmax attention Han et al. (2024). Another benefit of focusing on linear attention is that it allows us to isolate the role of the nonlinear MLP head in capturing the nonlinear nature of our ICL tasks since nonlinear attention can also play a role in capturing such nonlinear relations. In the rest of this section, we first describe the linear attention mechanism–introducing linear Transformer without MLP head as our baseline– and then, we explain our Transformer model with MLP head together with the training procedure.
@@ 3.3.1 Linear attention (linear Transformer)
where ${\bm{K}},{\bm{Q}},{\bm{V}}$ are appropriately sized key, query, and value matrices, respectively. When predicting $y_{\ell+1}$, the relevant output of the linear Transformer (without MLPs) will be $A_{d+1,\ell+1}$, i.e., the element of $\bm{A}$} corresponding to the 0 element in the embedding matrix $\bm{Z}$}. Using the reparameterization in Lu et al. (2024) (see Appendix A), this prediction can be simplified to:
@@ 3.3.2 Transformer with a nonlinear MLP
To simplify our analysis while keeping feature learning for the MLP, we consider the following training procedure with two stages Ba et al. (2022); Damian et al. (2022):
@@
Here, by avoiding data reuse, we ensure $\hat{{\bm{F}}}$ remains independent of the second training set, which simplifies theoretical analysis Ba et al. (2022). Furthermore, while our training procedure differs from standard end-to-end training techniques, this two-phase scheme –gradient descent on the first layer followed by ridge regression on the second– is well-motivated by prior theoretical work on MLPs Ba et al. (2022); Damian et al. (2022). This separation enables tractable analysis while preserving meaningful feature learning dynamics.
@@ 3.4 Evaluating ICL Performance
where $\hat{y}$ denotes the model prediction, given by (4) for the linear Transformer or by (5) for the Transformer with a nonlinear MLP. This formulation enables controlled experimentation with training-time data mixtures while evaluating performance as an average across all sources. Such an approach aligns with practical multi-source training scenarios, where models are trained on heterogeneous datasets and validated by averaging performance over the constituent sources Ye et al. (2025). Note that, we also provide ICL errors corresponding to each source (for cases when it might be relevant) in Appendix G.
@@ 4.1 Assumptions
This assumption aligns with regimes studied in prior work: $\ell/d$ and $n/d^{2}$ are critical for ICL performance in linear Transformers Lu et al. (2024), while $k/n$ ensures the model capacity scales with data Hu and Lu (2023).
@@
The covariance matrices of input vectors satisfy $\|{\bm{\Sigma}}_{x,s}\|_{2}^{2}=\mathcal{O}(d)$ and $\operatorname{Tr}({\bm{\Sigma}}_{x,s})\asymp d$ while the mean vectors satisfy $\|{\bm{\mu}}_{x,s}\|_{2}^{2}=\mathcal{O}(d^{1/2})$.
@@
By permitting step size to scale with the input dimension while the dimension tends to infinity, we introduce nontrivial feature learning, allowing the model to learn nonlinear features Ba et al. (2022); Demir and Dogan (2025a).
@@
$\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})|s=i))=\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})|s=j)),$
@@
$\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})|s)={\bm{I}}_{d(d+1)}+\sum_{q=1}^{r_{s}}\theta_{s,q}\bm{\gamma}_{s,q}\bm{\gamma}_{s,q}^{T},$ (9)
where $r_{s}\in\mathbb{Z}^{+}$, $\theta_{s,q}>0$, and $\{\bm{\gamma}_{s,q}\}$ are orthonormal vectors in $\mathbb{R}^{d(d+1)}$.
@@
$\eta\cdot\|\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))\|_{2}=o(d^{2}).$
Assumptions 4.3- 4.5 ensure tractability in our asymptotic analysis while allowing the feature learning. Although these assumptions might look unnatural at first glance, they enable us to connect our setting involving ICL by Transformer with a nonlinear MLP to prior results Demir and Dogan (2025a) for MLPs, facilitating our theoretical analysis. We would like to note that Assumptions 4.3- 4.5 represent limitations of our theoretical results. Nevertheless, our experimental results on both synthetic and real-world datasets suggest that these assumptions can be partially relaxed in practice.
@@
The vector ${\bm{w}}\sim\mathcal{N}(0,\frac{{\bm{I}}_{k}}{k})$ and the matrix ${\bm{F}}:=[{\bm{f}}_{1},\dots,{\bm{f}}_{k}]^{T}$ has rows ${\bm{f}}_{i}\sim\mathcal{N}\left(0,\frac{{\bm{I}}_{d(d+1)}}{\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))}\right)$.
This initialization is consistent with standard practices for MLP initializations Goodfellow et al. (2016).
@@
The activation $\sigma:\mathbb{R}\to\mathbb{R}$ has bounded derivatives and satisfies $\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma(z)^{2}]<\infty$, admitting a Hermite expansion (O’Donnell_2014, Chapter 11.2):